OpenStudy (anonymous):

Determine the electronegativity between the atoms of each molecule. Identify the bond as either ionic or covalent. State whether the molecule is polar or non polar. Identify the structure as having hydrogen bonding, dipole-dipole moments or London dispersion forces (LDF).

OpenStudy (anonymous):

i need these 4 questions for 4 elements: H2O, NH3, CH4, and CO2

OpenStudy (taramgrant0543664):

An ionic bond is formed when one atom accepts or donates one or more of its valence electrons to another atom. A covalent bond is formed when atoms share valence electrons. The atoms do not always share the electrons equally, so a polar covalent bond may be the result.

OpenStudy (anonymous):

so which element has covalent bonds?

OpenStudy (anonymous):

i think all of them

OpenStudy (anonymous):

It'll probably help to look at the Lewis structures of these compounds, as well as an electronegativity chart (all of which you can find online).

OpenStudy (taramgrant0543664):

The ones with covalent bonds are the ones that share their electrons

OpenStudy (anonymous):

so all of them?

OpenStudy (anonymous):

i already have all the lewis structures in front of me

OpenStudy (taramgrant0543664):

Yes all are covalent bonds of some sort

OpenStudy (anonymous):

ok, and im really having trouble with the electronegativity, how do i calculate that

OpenStudy (taramgrant0543664):

The electronegativity can be found on the periodic table and you subtract the lower one from the higher one

OpenStudy (anonymous):

so subtract the atomic number from atomic mass?

OpenStudy (taramgrant0543664):

No there is a Number on the back of the periodic table that is the electronegativity of the element you can also search it too

OpenStudy (anonymous):

and then subtract the electronegativity from the atomic number?

OpenStudy (taramgrant0543664):

No you subtract the electronegativities between the two elements

OpenStudy (anonymous):

im lost.. could you do an example? what two elements are you talking abuit

OpenStudy (taramgrant0543664):

HF H has an electronegativity of 2.20 F has an electronegativity of 3.98 3.98-2.20=1.78

OpenStudy (anonymous):

ohhh so you always subtract the lesser from the larger? or can you get a negative

OpenStudy (taramgrant0543664):

I don't think there should be a negative i'm pretty sure it is always larger minus smaller

OpenStudy (anonymous):

ok give me a sec to do all 4

OpenStudy (abhisar):

If \(\sf \Delta EN>1.7\), then the bond must be ionic. For deciding whether the molecule is polar or not, you need to consider the geometry of molecule. If it is symmetrical then its non-polar otherwise polar. Ih any molecule has a hydrogen atom attached to highly electronegative elements like N, O or F, then there are greater chances of hydrogen bonding.

OpenStudy (anonymous):

H2O: Polar CO2: Nonpolar CH4: Nonpolar NH3: Polar im pretty sure im correct

OpenStudy (abhisar):

Example, \(\sf NH_3\) EN of N is 3.04 EN of H is 2.2 \(\sf \Delta EN=3.04-2.2=0.84<1.7\) so the bonds will be covalent

OpenStudy (anonymous):

they're all covalent

OpenStudy (abhisar):



OpenStudy (anonymous):

EN of H2O is 1.24?

OpenStudy (abhisar):

EN is for a element.

OpenStudy (anonymous):

oh, i meant 1.24 is what i got after i subtactred them

OpenStudy (abhisar):

NH3 has a tetrahedral structure with a lone pair, thus it is unsymmetrical and hence has a dipole moment which makes it a polar molecule.

OpenStudy (anonymous):

i see

OpenStudy (abhisar):

For \(\sf H_2O\) there are 2 O-H bonds, you need to see if O-H bond is ionic or covalent. \(\sf \Delta EN = 3.44-2.2=?\)

OpenStudy (abhisar):

I need to leave now, I'll be back in a hour or so. I'll request @taramgrant0543664 to help you further c:

OpenStudy (taramgrant0543664):

Yes you were correct about which ones are polar and non polar ones

OpenStudy (anonymous):

ok, and i was correct about them all being covalent right

OpenStudy (taramgrant0543664):

Yes they are only polar and non-polar if they are also covalent

OpenStudy (anonymous):

and i did the calculations and this is what i got H2O: 1.24 CO2: 0.89 CH4: 0.35 NH3: 0.84